Minutes of the Trigger Operation Meeting (, 2007)
Attendees: Debbie Bard, Brandon Drummond, Al Eisner, Fang Fang, Kim Hojeong, Su Dong.
- Thusday owl shift - paged for damage in TSF. It first appeared to be the 'usual' occasional out-of-synch dammage, but a reconfigure didn't stop the problem in the next run. The odf log files reveal the problem - reframe errors and spurious interrupts show up, which is indicative of a firmware glitch, for example if the PEP timing signal is scrambled. Power cycle of the TSF crates fixed the problem.
- There's a 3-day ROD next week (mon-weds). We have no work planned in that time.
- In some high background runs, the M plot (top left) shows a larger peak than normal. This is just because of the background distributions in the detector.
- Still need to run L1TNtuple over the test cosmics runs made with the new GLT libraries, for validation.
- Sunday swing: the DQM noticed a hole in the L1 accepts per
node plot. It turned out that the node was running but it
was disconnected so that it did not appear in the JAS plots
nor in the hbook files.
Reconfiguring and starting a new run fixed the problem, but it's never happened before. Also, what do we do about the lumi and event count loss from this node? It'll be small, but will it affect anything?
Data Quality and Monitoring
- See here for discussion of anomolous run lumis from Run 5 stripcharts.
This page is created by Debbie Bard (firstname.lastname@example.org); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.